# Boosting Open Set Recognition Performance through Modulated Representation Learning

### v1

@@ 4.3 Performance on the Semantic Shift Benchmarks
Here, we report the performance on the SSBs for both the previous OSR methods and our scheme in Table 2. We find only three previous methods in literature that report results on the SSBs. We implement the methods by Chen et al. (2021) and Vaze et al. (2022) by including our TS. As LS has shown performance improvements on SSBs, we experiment both with and without uniform LS for the CE and SupCon losses. Here, we do not optimize performance for $\alpha$ and temperatures but use a fixed set of hyperparameters for consistency. By following the strategy mentioned in Section 4.1, we fix $(\tau^{+},\tau^{-})$ at $(0.3,0.1)$ for TS and $\tau=0.2$ for reporting with constant temperatures in the SupCon loss. For CE loss, with the base temperature being $1.0$, we set $(\tau^{+},\tau^{-})$ at $(0.5,2.0)$.
@@
We observe that if we include the proposed NegCosSch using either loss function, we get improved performances both for the closed set and open set problems over the corresponding baselines for all three datasets except for only the case of Aircraft on SupCon loss and including LS. Our NegCosSch provides performance boost for up to 1.60% of accuracy, up to 3.56%/2.99% of open set AUROC in the ‘Easy’/‘Hard’ splits and up to 4.32%/3.72% of OSCR. Moreover, our method can be used together with LS to further boost the performance in a few cases. This amount of performance boost comes without any additional computational cost.
@@
For cases where LS does not help the open set performance (for the Stanford cars -‘Hard’ split and Aircraft with CE loss), our NegCosSch outperforms the baselines. Overall for SSBs, the CE loss performs better than SupCon loss. Comparison of these methods with CosSch is listed in the Appendix. Moreover, our method outperforms the OSR methods by Chen et al. (2021) and Vaze et al. (2022) in most cases. We believe that our TS scheme in principle can improve other OSR methods, such as the method by Jia et al. (2024).
@@
Figure 4: Effect on performance improvement for our proposed scheduling over the baselines with varying number of training classes. Increasing the number of training classes tends to yield greater improvements in OSCR across all datasets, along with significant improvements in AUROC and accuracy, with the effect being most pronounced for CUB and least for FGVC-Aircraft. Error bars represent the standard deviation.
4.4 Proposed scheduling is more beneficial with higher number of training classes
To show the strength of our scheme with an increased number of training classes, we train models on the CE loss with $\{20\%,40\%,60\%,80\%\}$ of the randomly chosen training classes for the SSBs without changing the open set. In Figure 4, we plot the improvements (averaged across 5 trials) over the corresponding baseline. We define improvement of any metric $m$ as
$\text{improvement}=m[\text{NegCosSch}(\tau^{+},\tau^{-})]-\max_{\tau\in\mathcal{T}_{CE}}\{m[\text{Const}(\tau)]\}$
We observe an overall upward improvement trend for most cases in both ‘Easy’ and ‘Hard’ OSR splits in all three datasets as the number of training classes increases. The negative values in Figure 4(c) is due to the fact that we compare our method with the maximum scores of three baselines. With more training classes, the task becomes harder for the base model, which is observed by the performance decline. Nonetheless, our scheduling scheme gains higher improvement through TS.
4.5 Limitations
The open set performance depends on the nature of the unknown classes and their semantic similarity with the known classes. We find that the improvements over the baselines decrease, with a slight drop in closed set accuracy, when there is a small number of training classes, which occur in the CIFAR10, CIFAR+10, and CIFAR+50 benchmarks.
@@ 5 Related Works
Open Set Recognition: Since the introduction of OSR problem, it has received a significant interest in the research community. Most of the research attempts can be summarized into several common categories. The early ones are based on extreme value theory that modify the softmax function to model the open set samples (Vignotto and Engelke (2018); Bendale and Boult (2016)). Several works use generative models to generate synthetic unknown or counter-factual samples (Ge et al. (2017); Neal et al. (2018); Chen et al. (2021); Moon et al. (2022)), while other add auxiliary samples in the dataset with strategies, such as the sample mix-up (Chen et al. (2021); Xu et al. (2023); Bahavan et al. (2025); Li et al. (2024)), manifold mix-up (Zhou et al. (2021)), or by adding extra augmented samples (Wang et al. (2025); Jiang et al. (2023); Jia et al. (2024); Xu and Keuper ). Another group of research depends on training an additional model for a secondary objective function (Oza and Patel (2019); Sun et al. (2020); Perera et al. (2020); Yoshihashi et al. (2019); Zhang et al. (2020); Jia et al. (2024); Zhou et al. (2024)), for example training VAEs for computing the reconstruction loss hoping that it can improves the OSR score. However, training generative models or a secondary model is tough for larger benchmarks as it requires significant computation overhead.
Some methods either construct a different loss function (Chen et al. (2021, 2020a); Wang et al. (2022)) or add regularization to bound the open set risks (Zhou et al. (2021); Lu et al. (2022); Yang et al. (2024a)). For example, the method by Zhou et al. (2021) learns additional place-holders for the novel classes. Methods by Chen et al. (2021, 2020a) learn the reciprocal points of known classes signifying the representations of the ‘otherness’ corresponding to each class. These methods try to create additional empty regions in the representation space hoping that open set representations lie in those regions. The paper (Vaze et al. (2022)) shows that a well-trained closed set classifier is essential for good OSR. This has triggered research for better representation learning. For example, methods by Xu et al. (2023); Xu (2024); Li et al. (2024); Bahavan et al. (2025); Li et al. (2025) train models using the contrastive loss with regularization and heavy augmentations (Wang et al. (2025); Jiang et al. (2023); Jia et al. (2024)). The method by Wang et al. (2024) trains multiple experts for extracting diverse representations, and Yang et al. (2024b) proposes an open set self-learning framework, which adapts the model according to the test data assuming that it is available.
Temperature Scaling: Temperature scaling in the CE loss plays a crucial role in knowledge distillation (Hinton et al. (2015)), model calibration (Guo et al. (2017)) and so on. Since the contrastive loss has become popular for many tasks (Chen et al. (2020b); Khosla et al. (2020)), studies have aimed to understand its behavior. Works in Wang and Liu (2021); Kukleva et al. (2023); Zhang et al. (2021, 2022) explain the role of temperature in self-supervised contrastive learning. Recently, methods in Jeong et al. (2024); Qiu et al. (2024) utilize temperature cool-down methods in language models.

### v2

@@ 5.2 Performance of our Temperature Schedules across Various Loss Functions
Here, we report the performance on different OSR loss functions by including and without our NegCosSch in Table 2. We implement the ARPL loss, the CE baseline by Vaze et al. (2022) and the widely implemented SupCon loss in the recent OSR literature (Xu et al., 2023; Zhou et al., 2024a). As label smoothing (LS) has shown performance improvements, we experiment both with and without uniform LS for the CE and SupCon losses. Here, we do not optimize performance for the LS coefficient and temperatures but use a fixed set of hyperparameters for consistency.
@@
We observe that including the proposed schedules (P-NegCosSch or M-NegCosSch) in any OSR loss function improves performances both for the closed set and open set problems over the corresponding constant temperature baseline for all cases except for two, such as for Aircraft on SupCon loss including LS and P-NegCosSch, and for TinyImageNet on CE loss including LS and M-NegCosSch. Our NegCosSch provides performance boost for up to 1.87% of accuracy, up to 3.3%/3.1% of open set AUROC in the ‘Easy’/‘Hard’ splits and up to 4.4%/3.96% of OSCR. This amount of performance boost comes without any additional computational cost. Between our two schedules, the M-NegCosSch performs better in most cases than the periodic one.
@@
Moreover, our method can be used together with LS to further boost the performance in a few cases. Even for cases where LS does not help the open set performance (for the Stanford cars -‘Hard’ split and Aircraft with CE loss), our NegCosSch outperforms the baselines. Overall for SSBs, the CE loss performs better than SupCon loss. We believe that our scheme, in principle, can improve other OSR methods, such as the method by Jia et al. (2024).
@@
Figure 3: Effect on performance improvement for our proposed schedule over the baselines with varying number of training classes. Increasing the number of training classes tends to yield greater improvements in OSCR across all datasets, along with significant improvements in AUROC and accuracy, with the effect being most pronounced for CUB and least for FGVC-Aircraft. Error bars represent the standard deviations across trials with random training classes.
5.3 Proposed schedule is more beneficial with more training classes
To show the strength of our NegCosSch with increased number of training classes, we train models on CE loss with $\{20\%,40\%,60\%,80\%,100\%\}$ of the randomly chosen training classes for the SSBs without changing the open set. In Figure 3, we plot the improvement of a metric $m$ over the corresponding baseline, defined as: $\text{improvement}=m[\mathcal{T}_{\text{NegCosSch}}(\tau^{+},\tau^{-})]-\max_{\tau\in\mathbb{T}_{CE}}\{m[\mathcal{T}_{\text{Const}}(\tau)]\}$. We observe an overall upward improvement trend with more training classes in most cases for both the ‘Easy’ and ‘Hard’ OSR splits. The negative values in Figure 3(c) is due to the fact that we measure improvement over the maximum score of three baselines. With more training classes, the task becomes harder for the base model, which is observed by the performance decline. Nonetheless, our schedule gains higher improvement with more training classes.
@@ 6 Related Works
Open Set Recognition. Since the introduction of OSR problem, it has received a significant interest in the research community. Most of the research attempts can be summarized into several common categories, some of which are discussed in Section 1. Besides the use of generative models, input or mani-fold mix-up, other works add auxiliary samples with different strong augmentations for training models (Wang et al., 2025; Jiang et al., 2023; Jia et al., 2024; Xu & Keuper, ). Another huge group of research depends on training an additional model with a secondary objective function (Oza & Patel (2019); Sun et al. (2020); Perera et al. (2020); Yoshihashi et al. (2019); Zhang et al. (2020); Jia et al. (2024); Zhou et al. (2024a)). However, training generative models or a secondary model is tough for larger benchmarks as it requires significant computation overhead.
Another set of methods either construct a different loss function (Chen et al. (2021; 2020a); Wang et al. (2022)) or add regularization to bound the open set risks (Zhou et al. (2021); Lu et al. (2022); Yang et al. (2024a)). For example, the method by Zhou et al. (2021) learns additional place-holders for the novel classes. Methods by Chen et al. (2021; 2020a) learn the reciprocal points of known classes representing the ‘otherness’ corresponding to each class. These methods try to create additional empty regions in the representation space hoping that open set representations lie in those regions. The new baseline by Vaze et al. (2022) with well-trained closed set classifiers has triggered the OSR research for better representation learning schemes. For example, methods by Xu et al. (2023); Xu (2024); Li et al. (2024); Bahavan et al. (2025); Li et al. (2025) train models using the contrastive loss with regularization and heavy augmentations (Wang et al. (2025); Jiang et al. (2023); Jia et al. (2024)). The method by Wang et al. (2024) trains multiple experts for extracting diverse representations, and Yang et al. (2024b) proposes an open set self-learning framework, which adapts the model according to the test data assuming that it is available. Furthermore, several prior works have focused on developing fine-grained OSR methods (Lang et al., 2024; Bao et al., 2023; Sun et al., 2023).
Temperature Scaling. Temperature scaling in the CE loss plays a crucial role in knowledge distillation (Hinton et al. (2015)), model calibration (Guo et al. (2017)) and so on. Since the contrastive loss has become popular for many tasks (Chen et al. (2020b); Khosla et al. (2020)), studies have aimed to understand its behavior. Recently, methods by Jeong et al. (2024); Qiu et al. (2024) utilize temperature cool-down in language models.